OpenStudy (anonymous):

Differentiate the given funtion: f(x)= ln (x+1)/(x-1) I get -2/(x-1)^2

OpenStudy (anonymous):

amistre

OpenStudy (anonymous):

wait is it ln(x+1/x-1)?

OpenStudy (anonymous):

who wats to share google doc

OpenStudy (anonymous):

who has g mail

OpenStudy (amistre64):

ln[(x+1)/(x-1)] = ln(x+1) - ln(x+1)

OpenStudy (amistre64):

hi mary :)

OpenStudy (anonymous):

amistre answer me

OpenStudy (amistre64):

1/(x+1) - 1/(x-1) is good for the derivative

OpenStudy (anonymous):

do you have google account

OpenStudy (amistre64):

lotsa people have gmail...

OpenStudy (anonymous):

wht to share on goole doc

OpenStudy (amistre64):

no thanx mary :)

OpenStudy (anonymous):

i need help there

OpenStudy (anonymous):

whats yuor gmail

OpenStudy (amistre64):

tony031172@gmail is my spam catcher

OpenStudy (anonymous):

lol

OpenStudy (amistre64):

(x-1) - (x+1) -2 ----------- = ------ is what I get (x-1)(x+1) x^2-1

OpenStudy (anonymous):

\[\frac{1}{(-1+x) (1+x)}-\frac{\text{Log}[1+x]}{(-1+x)^2} \] \[\int\limits \left(\frac{1}{(-1+x) (1+x)}-\frac{\text{Log}[1+x]}{(-1+x)^2}\right) \, dx=\frac{\text{Log}[1+x]}{-1+x} \]

OpenStudy (anonymous):

amistre plz check your gmail

OpenStudy (anonymous):

sorry i started a new one. too much going on here

OpenStudy (amistre64):

lincoln assassination....

OpenStudy (anonymous):

yes

OpenStudy (amistre64):

differentiate is derivatives or integrals?

OpenStudy (anonymous):

amistrre join

OpenStudy (anonymous):

plz

OpenStudy (amistre64):

mary: im good with math; not english lit projects :)

OpenStudy (anonymous):

idc just go to google doc

OpenStudy (anonymous):

John Boothe was the culprit

OpenStudy (anonymous):

Very sorry about the mutiple postings. This whole business is not working out for me.

OpenStudy (amistre64):

Boothe was framed!! lol

OpenStudy (anonymous):

ok chague do you have gamail

OpenStudy (anonymous):

join my google doc

OpenStudy (anonymous):

vp andrew jackson wanted 187 Lincoln, so he could get all the cheese, Boothe obliged